# Low-Complexity MUSIC Algorithm: From Finite-Precision Perspective

### v2

@@ Abstract
The high computational complexity of the multiple signal classification (MUSIC) algorithm is mainly caused by the subspace decomposition and spectrum search, especially for frequent real-time applications or massive sensors. In this paper, we propose a low-complexity MUSIC algorithm from finite-precision arithmetic perspective. First, we analyze the computational bottlenecks of the classic low-complexity randomized unitary-based MUSIC (RU-MUSIC), formulating this computational issue as an inner product problem. Then, a mixed-precision method is introduced to address this problem. Specifically, this method partitions summations in inner products into blocks, where intra-block computations use low-precision arithmetic and inter-block sums use high-precision arithmetic. To further improve computational accuracy, we develop an adaptive-precision method that supports adaptive block sizes and multiple precision levels. Finally, simulation results show that the proposed finite-precision MUSIC design achieves direction-of-arrival (DOA) estimation performance similar to that using full-precision arithmetic while reducing more than 50% computational cost.
@@ I Introduction
Motivated by the above observations, in this paper, we propose a low-complexity MUSIC algorithm from finite-precision perspective. First, the computational bottlenecks of the randomized U-MUSIC (RU-MUSIC) are analyzed, and then this computational issue is formulated as an inner product problem. Further, we introduce a mixed-precision (MP) method to address this problem. Specifically, this method partitions summations in inner products into blocks, where intra-block computations are implemented by low-precision arithmetic and inter-block sums are calculated by high-precision arithmetic. Moreover, an adaptive-precision (AP) method is presented to enhance computational accuracy, supporting adaptive block sizes and multiple precisions. Finally, simulation results demonstrate that the proposed finite-precision MUSIC design can achieve a close DOA estimation performance to that of full-precision arithmetic, while it can reduce over 50% computational cost. Moreover, the notations of this paper is shown in Tab. I.
TABLE I: Notations
Variables Definitions Variables Definitions
$M$ Number of sensors $N$ Number of DoAs
$T$ Number of snapshots $F$ Number of discrete angles
$B$ Block size of MP method $u_{k}$ $k$-th precision
$p$ Number of block $\mathcal{G}_{k}$ $k$-th group of AP method
$\gamma$ Target accuracy $m_{k}$ Size of $\mathcal{G}_{k}$
@@ II-A Signal Model
denoting the steering vector of the $n$-th source with $\lambda$ representing the wavelength and $d=\lambda/2$ representing sensor spacing, $\mathsf{j}$ is imaginary unit, $\mathbf{s}\left(t\right)=\left[s_{1}(t),s_{2}(t),\cdots,s_{N}(t)\right]^{\mathsf{T}}\in\mathbb{C}^{N\times 1}$ is the incoherent signals at time instance $t$, ${\bf n}(t)\sim\mathcal{CN}({\bf 0},\sigma_{n}^{2}{\bf I}_{M})$ is the complex additive white Gaussian noise (AWGN), and $\sigma_{n}^{2}$ is the variance of noise. Then, the covariance matrix of $\mathbf{x}\left(t\right)$ is
@@ 6             $y_{k}=y_{k}+b_{i}c_{i}$ in precision $u_{l}$
Figure 1: The illustration of the MP method. Using the computation of inner products $y={\bf b}^{\mathsf{T}}{\bf c}$ as an example, where $p=\lceil M/B\rceil$.
Notably, directly applying finite-precision arithmetic, especially low-precision arithmetic, for inner products introduces substantial computational errors [16]. To mitigate these errors, MP methods have been investigated in numerical linear algebra [17] and signal processing [13], which is summarized in Algorithm 2. Specifically, this method involves partitioning summations in inner products into blocks with block size $B$, where intra-block computations use low precision $u_{l}$ and then inter-block sums in high precision $u_{h}$, as depicted in Fig. 1. Here, we use it for efficient implementation of the MUSIC algorithm. Nevertheless, the MP method is restricted to artificially fixed block sizes and two precision levels, and cannot support adaptive block sizes or multiple precisions. To this end, we present a generalized AP method in the following subsection.
@@ 5             $y_{k}=y_{k}+b_{i}c_{i}$ in precision $u_{k}$
Remark 4 (Comparison the MP and AP Method).
Comparing Algorithm 2 and 3, it can be observed that the AP method adapts the block size based on a target accuracy and supports multiple precision levels. Moreover, the AP method provides higher accuracy but incurs greater computational cost compared to the MP method, as intra-block computations in the MP method are always performed using low-precision arithmetic.
@@ IV-2 Parameters of DOA estimation
As shown in Fig. 2, we evaluate the performance and computational cost of seven algorithms: MUSIC, U-MUSIC, root-MUSIC, RU-MUSIC, RU-MUSIC with pure low-precision, MP-based RU-MUSIC, and AP-based RU-MUSIC. First, Fig. 2a illustrates the estimated spectrum of these algorithms in a single trial at SNR = 20 dB. Notably, the AP-based RU-MUSIC achieves a spectral approximation that closely aligns with U-MUSIC and RU-MUSIC, outperforming both the MP-based variant and pure low-precision implementations. Then, to quantify the computational cost, Fig. 2b compares the number of additions and multiplications across different methods. Specifically, compared with full-precision RU-MUSIC ($\mathtt{fp64}$), the AP-based method reduces additions and multiplications by $55.32\%$ and $62.08\%$, respectively. At the same time, it incurs only marginal overhead compared to the MP method. Finally, Fig. 2c analyzes RMSE performance. At low SNR levels, low-precision arithmetic ($\mathtt{fp16}$) achieves accuracy comparable to full-precision baselines. At high SNR levels, AP-based RU-MUSIC maintains near full-precision RMSE performance, whereas low-precision and MP-based implementations exhibit significant degradation. This confirms the AP method’s superior performance against rounding errors. Collectively, AP-based RU-MUSIC offers a favorable balance between computational efficiency and estimation accuracy. Moreover, when computing resources are severely limited, MP-based RU-MUSIC is a good choice.

### v4

# Revisiting MUSIC: A Finite-Precision Perspective

@@ Abstract
The high computational complexity of the multiple signal classification (MUSIC) algorithm is mainly caused by the subspace decomposition and spectrum search, especially for frequent real-time applications or massive sensors. In this paper, we propose a low-complexity MUSIC algorithm from a finite-precision arithmetic perspective. First, we analyze the computational bottlenecks of the classic low-complexity randomized unitary-based MUSIC (RU-MUSIC), formulating this computational issue as an inner product problem. Then, a mixed-precision method is introduced to address this problem. Specifically, this method partitions summations in inner products into blocks, where intra-block computations use low-precision arithmetic and inter-block sums use high-precision arithmetic. To further improve computational accuracy, we develop an adaptive-precision method that supports adaptive block sizes and multiple precision levels. Finally, simulation results show that the proposed finite-precision MUSIC design achieves direction-of-arrival (DOA) estimation performance similar to that using full-precision arithmetic while reducing more than 50% computational cost.
@@ I Introduction
Motivated by the above observations, in this paper, we propose a low-complexity MUSIC algorithm from a finite-precision perspective. First, the computational bottlenecks of the randomized U-MUSIC (RU-MUSIC) are analyzed, and then this computational issue is formulated as an inner product problem. Further, we introduce a mixed-precision (MP) method to address this problem. Specifically, this method partitions summations in inner products into blocks, where intra-block computations are implemented by low-precision arithmetic and inter-block sums are calculated by high-precision arithmetic. Moreover, an adaptive-precision (AP) method is presented to enhance computational accuracy, supporting adaptive block sizes and multiple precisions. Finally, simulation results demonstrate that the proposed finite-precision MUSIC design can achieve a close DOA estimation performance to that of full-precision arithmetic, while it can reduce over 50% computational cost.
@@ II-A Signal Model
denoting the steering vector of the $n$-th source with $\lambda$ representing the wavelength and $d=\lambda/2$ representing sensor spacing, $\mathsf{j}$ is imaginary unit, $\mathbf{s}\left(t\right)=\left[s_{1}(t),s_{2}(t),\cdots,s_{N}(t)\right]^{\mathsf{T}}\in\mathbb{C}^{N\times 1}$ is the incoherent signals at time instance $t$, ${\bf n}(t)\sim\mathcal{CN}({\bf 0},\sigma_{n}^{2}{\bf I}_{M})$ is the complex additive white Gaussian noise (AWGN), $\sigma_{n}^{2}$ is the variance of noise, and $T$ is the number of snapshots. Then, the covariance matrix of $\mathbf{x}\left(t\right)$ is
@@ 6             $y_{k}=y_{k}+b_{i}c_{i}$ in precision $u_{l}$
Notably, directly applying finite-precision arithmetic, especially low-precision arithmetic, for inner products introduces substantial computational errors [16]. To mitigate these errors, MP methods have been investigated in numerical linear algebra [17] and signal processing [13], which is summarized in Algorithm 2. Specifically, this method involves partitioning summations in inner products into blocks with block size $B$, where intra-block computations use low precision $u_{l}$ and then inter-block sums in high precision $u_{h}$. Here, we use it for efficient implementation of the MUSIC algorithm. Nevertheless, the MP method is restricted to artificially fixed block sizes and two precision levels, and cannot support adaptive block sizes or multiple precisions. To this end, we present a generalized AP method in the following subsection.
@@ IV-2 Parameters of DOA estimation
As shown in Fig. 1, we evaluate the performance and computational cost of seven algorithms: MUSIC, U-MUSIC, root-MUSIC, RU-MUSIC, RU-MUSIC with pure low-precision, MP-based RU-MUSIC, and AP-based RU-MUSIC. First, Fig. 1a illustrates the estimated spectrum of these algorithms in a single trial at SNR = 20 dB. Notably, the AP-based RU-MUSIC achieves a spectral approximation that closely aligns with U-MUSIC and RU-MUSIC, outperforming both the MP-based variant and pure low-precision implementations. Then, to quantify the computational cost, Fig. 1b compares the number of additions and multiplications across different methods. Specifically, compared with full-precision RU-MUSIC ($\mathtt{fp64}$), the AP-based method reduces additions and multiplications by $55.32\%$ and $62.08\%$, respectively. At the same time, it incurs only marginal overhead compared to the MP method. Finally, Fig. 1c analyzes RMSE performance. At low SNR levels, low-precision arithmetic ($\mathtt{fp16}$) achieves accuracy comparable to full-precision baselines. At high SNR levels, AP-based RU-MUSIC maintains near full-precision RMSE performance, whereas low-precision and MP-based implementations exhibit significant degradation. This confirms the AP method’s superior performance against rounding errors. Collectively, AP-based RU-MUSIC offers a favorable balance between computational efficiency and estimation accuracy. When computing resources are severely limited, MP-based RU-MUSIC is a good choice.